x²y + 5. Find the exact value of lim (5,7,2)=(0,0,0) 3&4 + 4y2 + 2yz? if it exists. 9 -524
Answer #1

lim_{(x,y,z)\rightarrow (0,0,0)} \frac{x^2y+2yz^2}{3z^4 + 4y^2 + 5z^4} = lim_{h\rightarrow 0} \frac{h^2.h + 2h.h^2}{3h^4 + 4h^2 + 5h^4} = lim_{h\rightarrow 0} \frac{3h^3}{4h^2 + 8h^4} = lim_{h\rightarrow 0} \frac{3h}{4 + 8h^2} = 0

We have to cancel h^2 from both numerator and denominator in order to avoid 0/0, now it is in 0/finite form, so the exact value will be 0

Note - Post any doubts/queries in comments section.